Hello. Can I get the answer for D, E and F and please give a clear explanation on F. Thank you
CHAPTER 21 LEASING Question 1 Your firm is considering leasing a new computer. The lease lasts for 9 years. The lease calls for 10 payments of $1,000 per year with the first payment occurring immediately. The computer would cost $7,650 to buy and would be straight-line depreciated to a zero-salvage value over 9 years. The actual salvage value is negligible because of technological obsolescence. The firm carn borrow at a rate of 8%. The corporate tax rate is 30%. a) Calculate the appropriate discount rate for valuing the lease (b) Calculate the after-tax cash flow from leasing relative to the purchase in years 1-9. (e) Calculate the after-tax cash flow from leasing relative to the purchase in year 0 d Calculate the NPV of the lease relative to the purchase. e) Calculate the after-tax cash flow in year 9 be if the asset had a residual value of $500 oDoes this type of lease is capital lease or operating lease? Explain
Answer #1

D:

NPV = Initial cost+ PV of cash flows

We calculate the incremental cash flows from leasing

Year 0= 7650

Years 1 to 9 = Lease payments -Tax shield on investment

= -1000-30%*7650/9

=-1255

NPV = 7650- 1255*(1-1/1.08^9)/0.08

= -189.84

E:INCREMENTAL After tax cash flow in year9 = Incremental lease payments – residual value*(1-tax)

= -1255- 500*(1-0.3)

= =1605

F: This is an operating lease since it is treated like rent payments. The asset is not purchased in this case. In a capital lease the payments are treated like loan repayments.